OpenStudy (anonymous):

Study session.... emergency! please help?? What is a possible first step in eliminating a variable in the following system of equations?x + 7y = -15 2x + 7y = -9 Answer Multiply the first equation by -1. Multiply the second equation by -1 Multiply the first equation by -2. All of the above

OpenStudy (anonymous):

(I think it's C...)

OpenStudy (anonymous):

All of them..

OpenStudy (anonymous):

@Kashan Thank you...........!!!!!

OpenStudy (anonymous):

all of them

OpenStudy (anonymous):

@InsanelyChaotic no problem... :)

OpenStudy (anonymous):

A) x+7-7y=-15 first subtract 7 from both side. then it will become x-7y=-22 - 2x+7y=-9 the -7y and +7y are gone. so =-x=-13 x=13 B) x+7-7y=-15 add 7 to both sides. = x-7y=-22 (-) x-7y=-22 = -31 13, -31 i think thats is the answer. I HOPE ITS HELPFUL! :D